2)solve (2x^2+2xy+2xz^2+1)dx+dy+2zdz=0


1
Expert's answer
2021-06-04T03:49:08-0400

"P=2x^\n2\n +2xy+2xz^\n2+1\n ,Q=1,R=2z"

If x is a constant ", dx=0,"

"dy+2zdz=0,"

"\int dy+\int 2zdz=C,"

"y+z^2=f(x)."

Differentiate "y+z^2=f(x),"

"dy+2zdz-f'(x)dx=0," "f \n'\n (x)=\u2212(2x^\n2\n +2xy+2xz^2+1)," "f'(x) =-(2x^2+1) - 2x( y+z^2\n )."

Replace "y+z^2=f(x),"

"f'(x) =-(2x^2+1)-2xf(x)",

"f \n'\n (x)+2xf(x)=(\u22122x ^\n2\n \u22121)".

Solve the linear equation "f'(x)+2xf(x)=\u2212(2x^\n2\n +1),"

"e^{\n \u222b2xdx}=e^{\n x^\n2\n},"

"e^{x^2}f(x)=-\int e^{x^2}(2x^2+1)dx+C="

"-\int x\cdot2xe^{x^2}dx-\int e^{x^2}dx+C="

"-xe^{x^2}+\int e^{x ^2}dx- \int e^{x^2} dx+C="

"=\u2212xe ^\n{\n x ^\n2\n }+C,"

"f(x) =-x+Ce^{-x^2\n },"

"y+z^2=-x+Ce^{-x^2},"

"x+y+z^2=Ce ^\n{\n \u2212x ^\n2\n }"